# LCMDC: Large-scale Chinese Medical Dialogue Corpora for Automatic Triage and Medical Consultation

### v1

@@ 5.4.1 Overall Performance
Table 5: Comparison of Text Generation Results
@@
BLEU-1 $\uparrow$ 0.0922 0.1214 0.1211
CHRF $\uparrow$ 0.0374 0.0769 0.0781
GLEU $\uparrow$ 0.0448 0.0977 0.1029
NIST $\uparrow$ 0.5398 1.3932 1.4049
RIBES $\uparrow$ 0.0637 0.0827 0.0903
TER $\downarrow$ 1.3519 1.0489 0.9997
WMD $\uparrow$ 0.5168 0.6192 0.6461
BERT F1 $\uparrow$ 0.568 0.646 0.648
@@
According to Table 5, the proposed model outperforms the traditional deep learning sequence-to-sequence (seq2seq) algorithm across almost all evaluation metrics. TER (Translation Edit Rate) measures the number of edits required to change a generated sentence to match the reference sentence, indicating the proximity between the generated and real texts. Therefore, a lower TER value implies a closer match between the generated and real texts. The superior performance of the model on these evaluation metrics is attributed to the knowledge embedded in the large model and the rationality of the proposed framework.
@@
Ablation study results are presented in Appendix D.2
6 Discussion and Future Work
We discuss the biases in online medical data, AI adoption in healthcare, and the development of large language models in different sectors. For details, please refer to Appendix E.
Looking ahead, future work focuses on expanding datasets to reduce bias, enhancing intelligent triage systems with background knowledge and multi-modal methods, and refining medical consultation systems through alignment between LLM outputs and human preferences. The potential for further applications of LLMs is vast, enabling more flexible and realistic solutions in various fields. For details, please refer to the Appendix F.
7 Conclusion
We collected and organized Chinese medical data, including datasets for automated medical triage and consultation, forming the foundation of our work. Based on these datasets, we developed two systems: an intelligent triage system using text classification and a medical consultation system using a generative language model. Experimental results and ablation studies validated the effectiveness of both systems, demonstrating improved performance in medical diagnosis and consultation tasks.
8 Limitation
Dataset Biases:
  • Lack of Emergency Data: The dataset lacks emergency and severe cases, limiting the model’s ability to generalize to critical or urgent medical situations.
  • Imbalanced Age Distribution: The dataset is biased towards younger users, with insufficient data on elderly populations, reducing the model’s applicability to older age groups.
  • Issues Due to Anonymity: The anonymity of online consultations often leads to more private or informal questions, which may not accurately reflect typical clinical scenarios, affecting the model’s real-world performance.
Intelligent Triage System:
  • Lack of Comparison with Human Performance: Current classification methods lack sufficient comparison with professional doctors’ diagnostic results, making assessing their reliability in practical medical applications difficult.
  • Lack of Multi-modal Research: Existing classification methods are primarily text-based and do not incorporate multi-modal data such as images or audio, limiting the understanding of comprehensive patient information.
Medical Consultation System:
  • No Multi-turn Dialogues: Generation methods are often limited to single-turn dialogues, lacking the ability to handle contextual information across multiple rounds of interaction, which is crucial in real consultations.
  • No Role-playing or Simulation: There is a lack of role-playing or simulation of real medical scenarios in the generated dialogues, resulting in responses that lack authenticity and practical relevance to clinical settings.
9 Ethical Issues
Our study does not present any ethical issues that would require formal ethics approval, as it does not involve human subjects, sensitive personal data, or any other ethical considerations that would necessitate formal ethics oversight. First, the anonymized patient-doctor dialogues were legally and compliantly obtained from publicly available sources, specifically from an online medical consultation platform. Second, users have agreed to the platform’s terms of service, which include being informed that these anonymized consultation dialogues may be made public, so the data has been made available to the public with the informed consent of the users.
Appendix A Related Work
A.1 Dataset
CBLUE Zhang et al. (2022) is a comprehensive Chinese biomedical language understanding benchmark designed to advance research in biomedical NLP for the Chinese language. MedDG Liu et al. (2020) is a large-scale dataset aimed at developing conversational agents for primary clinical advice, especially valuable during the COVID-19 pandemic. MedDialog Zeng et al. (2020) is a comprehensive dataset for training and evaluating medical dialogue systems, enhancing remote healthcare services. MedQuAD Ben Abacha and Demner-Fushman (2019) is a significant resource for developing AI models in medical question answering, providing a robust foundation for understanding and responding to medical inquiries. There are some other datasets like event recognition Deng et al. (2022).
A.2 Triage System
IntelTriage Billis et al. (2019) is an intelligent triage system for Emergency Departments in Greece, dynamically prioritizing patients and monitoring their vital signs using wearable biosensors. The Babylon AI-powered Triage and Diagnostic System Razzaki et al. (2018) demonstrated diagnostic accuracy comparable to human doctors and provided generally safer triage recommendations. MESTRIMAN Sierra et al. (1993) is an expert system designed to support medical assistance during catastrophes by coordinating resources and enhancing disaster management. The SADE model Li et al. (2024) addresses challenges in medical triage and diagnosis by using dual encoders to represent patient consultations and doctor diagnoses.
A.3 Question Answer System
This research Abdallah et al. (2020) automates the generation of qualified medical answers using an end-to-end RNN and encoder-decoder model trained on data from online health services. Another system Jiang et al. (2021) integrates medical knowledge, knowledge graphs, and QA systems, using crawler technology and rule-based algorithms for classification and querying. An end-to-end character-level multi-scale CNN framework Zhang et al. (2017) is introduced for Chinese medical question-answer matching. Med-PaLM 2 Singhal et al. (2023) leverages large language model improvements and medical domain finetuning to enhance performance on medical QA datasets.
Appendix B Dataset Details
B.1 Reasons for Data Source
Here are several reasons for choosing "Quick Doctor" website:
Appendix C Experiment Setup Details
C.1 Datasets
In Section 2, we collect medical corpus and formulate 2 datasets for the triage system and consultation system respectively. In the intelligent triage dataset, the inputs are the users’ questions, and the outputs are two different grained labels, indicating high-level and detailed triage.
C.2 Baselines
For the triage system, we employ SVM Cortes and Vapnik (1995) and decision tree Quinlan (1986) as the mechanism learning baselines, while TextCNN Kim (2014), RCNN Lai et al. (2015), LSTM Hochreiter and Schmidhuber (1997), Self-Attention Vaswani et al. (2017), and Capsule Network Sabour et al. (2017) are used as neural network baselines. Then We also utilize pre-trained language models including BERT, RoBERTa Liu et al. (2019), and MacBERT Cui et al. (2020). For the consultation system, we employ the Seq2Seq method and BART Lewis et al. (2020) as baselines.
C.3 Evaluation Metrics
For the triage system, we utilize classification metrics, including accuracy, precision, recall, and F-1 score. For the consultation system, we employ various evaluations, including n-gram overlap measures such as BLEU Papineni et al. (2002), NIST Doddington (2002), RIBES Isozaki et al. (2010), CHRF Popović (2015), and GLEU Mutton et al. (2007); distance-based measures like TER Snover et al. (2006); vector similarity measures including WMD Kusner et al. (2015); n-gram diversity measures such as Self BLEU; sentence semantic similarity measures including BERT score Zhang et al. (2019) and Web Bert Similarity; and other metrics such as Lexical Diversity, and KL Divergence.
C.4 Hyper Parameters
For continuous pre-training of BERT and MacBERT, we set the max input length as 128 and mask 15% of tokens in the MLM task. For the supervised training of the triage system, we set the embedding dimension to 768, and the LSTM hidden size to 1024. We train 50 epochs. The learning rate for the pre-trained language model is 5e-5, while that for other parts of the whole framework is 2e-4. For the prompt learning method, we train 20 epochs with a learning rate of 2e-5. For the consultation system, we use GPT-2 as the base model. We set the token embedding dimension to 768, and the vocabulary length of natural language tokens to 50257. In the pre-training stage, we trained 10 epochs using background knowledge. In the fine-tuning stage, we trained 20 epochs using a learning rate of 2.6e-5.
C.5 Experimental Environment
Appendix D Ablation Study Details
D.1 Supervised Triage System Experiment
To verify the necessity of each component in the supervised triage system , we conducted ablation experiments by removing each component.
(a) Coarse-grained
(b) Fine-grained
Figure 7: Ablation Results of Supervised Classification Method.
D.2 Medical Consultation System Experiment
@@
Table 6: Results of Text Generation Ablation Experiments
@@
The results in Table 6 demonstrate that each step in the proposed framework is crucial. Removing any step results in a decline in the model’s performance across various evaluation metrics, highlighting the superiority and rationality of the proposed structure.
@@
Appendix E Discussion
E.1 Data Issues
During data collection, we found certain biases in the data gathered from online platforms. Firstly, acute and severe cases are not typically presented online, as patients in these situations usually go directly to the hospital. Secondly, the internet acts as a filter, with the online population differing from the general population. This results in a majority of younger users seeking online consultations, leading to a lack of data on elderly conditions. Lastly, due to the anonymity of the internet, the nature of questions asked online can differ from those asked offline, often including more private issues.
Medical issues require maximum caution, and licensed doctors undergo extensive training to earn their qualifications due to the severe consequences of medical errors, which can endanger lives. Therefore, using current AI, which often functions as a black-box system, for diagnosis and treatment remains an ethical and legal grey area. At present, AI can only serve as a supplementary tool. Similar to autonomous driving, where accidents pose challenges in accountability, AI in healthcare must develop gradually. Automation systems must advance in a phased manner.
E.2 Discussion on Current Large Models
With the rise of ChatGPT, pre-trained language models seem to have deviated from their original purpose but have unified various tasks. Previously proposed task classifications are now outdated. Since the introduction of the T5 model, all-natural language processing tasks can be modeled as text generation tasks, differing only in prompts. Currently, the boundaries between various big data tasks are unclear, and many tasks can be converted into generative tasks. The popularity of ChatGPT has even given rise to a new role, "prompt engineer," who specializes in crafting effective prompts to leverage ChatGPT for task completion. In the future, large models can liberate and enhance productivity, handling repetitive tasks and freeing humans for creative work, thus advancing technological progress.
E.3 Comparison Between Industrial and Research Institutions in Large Model Development
Industries aim to generate profits, driving them to develop market-appropriate large models with high enthusiasm. They also have vast user bases that provide data and accelerate training. However, industries face ethical and legal constraints, such as patents and privacy issues. Research institutions, lacking immediate application needs or access to first-hand data, face disadvantages in a data-driven era. Nonetheless, they pursue innovation without profit motives, experimenting in numerous directions, where even a one-in-a-million success marks a significant breakthrough. Furthermore, research institutions frequently refresh their teams, maintaining high efficiency in research.
Appendix F Potential Future Work
F.1 Dataset Extension
Collecting data from multiple sources can help mitigate the bias faced in the dataset.
F.2 Intelligent Triage
The All-in-one LLMs can solve the triage in a more general way. Researchers may pay more attention to adding more background information, such as multi-modality methods. Double checking will ground truth after the decision is also important due to the high risk of medical problems.
F.3 Medical Consultant
Medical consultation systems have a promising future as AI technology advances, with human resources and potentially unlimited computing resources that can scale with demand. With the development of current LLMs, constructing a consultation system become easier. Researchers can shift their focus to the alignment between LLM output and human preference (legal issues, correct knowledge, cost trade-off). The RLHF may contribute to the alignment.
F.4 Further Application
The power of LLMs may empower more complicated tasks. As a result, researchers may not need to formulate real tasks into fixed mathematical tasks, which will enable more flexible and realistic solutions. Researchers may extend the usage of our datasets based on their own need.

### v2

@@ VI-D1 Overall Performance
TABLE V: Comparison of Text Generation Results
@@
BLEU-1 [49] $\uparrow$ 0.0922 0.1214 0.1211
CHRF [50] $\uparrow$ 0.0374 0.0769 0.0781
GLEU [51] $\uparrow$ 0.0448 0.0977 0.1029
NIST [52] $\uparrow$ 0.5398 1.3932 1.4049
RIBES [53] $\uparrow$ 0.0637 0.0827 0.0903
TER [54] $\downarrow$ 1.3519 1.0489 0.9997
WMD [55] $\uparrow$ 0.5168 0.6192 0.6461
BERT F1 [56] $\uparrow$ 0.568 0.646 0.648
@@
According to Table V, the proposed model outperforms the traditional deep learning sequence-to-sequence (seq2seq) algorithm across almost all evaluation metrics. TER (Translation Edit Rate) measures the number of edits required to change a generated sentence to match the reference sentence, indicating the proximity between the generated and real texts. Therefore, a lower TER value implies a closer match between the generated and real texts. The superior performance of the model on these evaluation metrics is attributed to the knowledge embedded in the large model and the rationality of the proposed framework.
@@
VI-D2 Ablation study
@@
TABLE VI: Results of Text Generation Ablation Experiments
@@
The results in Table VI demonstrate that each step in the proposed framework is crucial. Removing any step results in a decline in the model’s performance across various evaluation metrics, highlighting the superiority and rationality of the proposed structure.
@@
VII Conclusion
We collected and organized Chinese medical data, including datasets for automated medical triage and consultation, forming the foundation of our work. Based on these datasets, we developed two systems: an intelligent triage system using text classification and a medical consultation system using a generative language model. Experimental results validated the effectiveness of both systems, demonstrating improved performance in medical diagnosis and consultation tasks.
Looking ahead, future work focuses on expanding datasets to reduce bias, enhancing intelligent triage systems with background knowledge and multi-modal methods, and refining medical consultation systems through alignment between LLM outputs and human preferences. At the same time, inspired by the field of recommendation systems [57, 58], further considering rich user information combined with the powerful generalization ability of LLM to achieve personalized triage is also a future direction. The potential for further applications of LLMs is vast, enabling more flexible and realistic solutions in various fields.